The demand equation for a product is given by
P=30-0.1q2 .
(a)At what point price is demand inelastic?
(b)at what price is demand unitary elastic
1
Expert's answer
2018-02-15T13:56:08-0500
P’ = -0.2q
dQ/dP = -1/0.2q

(a) E = (dQ/dP)*(P/Q)
p/0.2q^2 = 0
p=0 or Q = 0.

(b) E = (dQ/dP)*(P/Q) = 1
p/0.2q^2 = 1
p = 0.2q^2
For example:
p = 20. Q = 10